# Interpretability by Design for Efficient Multi-Objective Reinforcement Learning

### v1

@@ Abstract
Multi-objective reinforcement learning (MORL) aims at optimising several, often conflicting goals in order to improve flexibility and reliability of RL in practical tasks. This can be achieved by finding diverse policies that are optimal for some objective preferences and non-dominated by optimal policies for other preferences so that they form a Pareto front in the multi-objective performance space. The relation between the multi-objective performance space and the parameter space that represents the policies is generally non-unique. Using a training scheme that is based on a locally linear map between the parameter space and the performance space, we show that an approximate Pareto front can provide an interpretation of the current parameter vectors in terms of the objectives which enables an effective search within contiguous solution domains. Experiments are conducted with and without retraining across different domains, and the comparison with previous methods demonstrates the efficiency of our approach.
@@ 1 Introduction
While a globally continuous mapping is an ideal, we notice that in non-trivial problems, the relationship between the performance space and parameter space of policies is not a simple, single continuous mapping but can be described by a family of locally continuous components (Xu et al., 2020; Li et al., 2024). Our findings suggest that effectively exploring just a few of these components can be sufficient to achieve competitive performance in typical benchmark problems. This understanding forms the basis of our core concept: the Parameter-Performance Relationship (PPR). The process is seeded by an RL task that finds a good but not necessarily optimal parameter vector for an initial policy. Then a second policy is obtained by retraining with different preferences, and from there, additional policies are efficiently generated by a locally linear extrapolation, which led to the name LLE-MORL for the approach that we present in the following. If the policies obtained by the extension process are briefly retrained, they can improve with further extension, although eventually they may become dominated by earlier solutions which would indicate the need for a restart with a different initial policy. Within each solution component, the policy representation is easily interpretable in terms of the continuous PPR, but also the boundaries where the policies depart from optimality are interesting. They indicate that a discontinuous reparametrisation takes place and that thus policies of a potentially qualitatively different type are optimal on either side of the boundary.
Building upon the notions of PPR and locally linear extension, in this paper, we introduce LLE-MORL, a MORL algorithm that is designed to efficiently trace the Pareto front (see Sect. 2.2) by systematically exploring these identified local structures. Our experiments demonstrate that the proposed algorithm can achieve high-quality Pareto front approximations with notable sample efficiency. This strong performance is primarily attributed to its simple yet effective locally linear extension method, which significantly reduces the need for extensive retraining along the Pareto front. Such efficiency is made by exploiting the locally continuous nature of the parameter-performance relationship, a characteristic that also enhances the overall interpretability of our approach.
@@ 2.1 Multi-Objective Reinforcement Learning
Definition 1
Multi-Objective Markov Decision Process (MOMDP). A MOMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\{\mathcal{R}^{d}\},\gamma,\Omega,f_{\Omega})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}(s^{\prime}|s,a)$ is the state transition probability, $\mathcal{R}^{d}$ is a vector-valued reward function with $d$ as the number of objectives, specifying the immediate reward for each of the considered objectives, $\gamma$ is the discount factor, $\Omega$ is the preferences space, $f_{\Omega}:\mathbb{R}^{d}\to\mathbb{R}$ is the scalarisation function.
The crucial difference between MOMDPs and traditional single-objective MDPs is the reward structure. While single-objective MDPs use a scalar reward function $\mathcal{R}$, MOMDPs feature a vector-valued reward function $\mathcal{R}^{d}$ that delivers distinct numeric feedback for each objective, directly correlating the length of the reward vector with the number of objectives. At each timestep $t$, the agent in state $s_{t}\in\mathcal{S}$ selects an action $a_{t}\sim\pi(\cdot\mid s_{t})$, transitions to a new state $s_{t+1}$ with probability $P(s_{t+1}\mid s_{t},a_{t})$, and receives a reward vector $\mathbf{r}_{t}=\bigl{[}(R_{1}(s_{t},a_{t}),\,R_{2}(s_{t},a_{t}),\,\dots,\,R_{d}(s_{t},a_{t})\bigr{]}).$ We define the discounted return vector by $\mathbf{G}_{t}=\sum_{k=0}^{\infty}\gamma^{k}\,\mathbf{r}_{t+k},$ and the multi-objective action-value function of a policy $\pi$ for a given state-action pair $(s,a)$ by $\mathbf{Q}^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl{[}\mathbf{G}_{t}\mid s_{t}=s,\,a_{t}=a\bigr{]}$. The goal of MORL is to find a policy $\pi$ such that the expected return of each objective can be optimised. In practice, we trade off objectives via a scalarisation function $f_{\omega}(\mathbf{r})$, which produces a scalar utility using preference vector $\omega\in\Omega$. The scalarisation function $f_{\omega}(\mathbf{r})$ is used for mapping the multi-objective reward vector $\mathbf{r}(s,a)$ to a single scalar. In this paper, we consider the linear scalarisation function $f_{\omega}(\mathbf{r}(s,a))=\mathbf{\omega}^{\mathbf{T}}\mathbf{r}(s,a)$, which is commonly used in MORL literature (Yang et al., 2019; Felten et al., 2024). When the preference dimension $d=1$ (so that the return vector is one-dimensional), the MOMDP collapses to a standard single-objective MDP, since the reward vector reduces to a scalar and $f_{\Omega}$ becomes the identity mapping.
@@ 2.2 Pareto Optimality
Definition 2
Pareto optimality A policy $\pi$ is said to dominate another policy $\pi^{\prime}$ if and only if: $\forall i\in\{1,\dots,d\},\quad V_{i}^{\pi}(s)\geq V_{i}^{\pi^{\prime}}(s),\quad\text{and}\quad\exists j,\,V_{j}^{\pi}(s)>V_{j}^{\pi^{\prime}}(s),$ where $V_{i}^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t})\mid s_{0}=s]$ denotes the expected discounted return for objective $i$ under policy $\pi$. A policy $\pi^{*}$ is Pareto optimal if and only if it is not dominated by another policy. The set of all Pareto optimal policies forms the Pareto set: $\mathcal{P}=\{\pi\mid\pi\text{ is Pareto optimal}\}.$ The corresponding set of expected returns incured by policies in the Pareto set is termed Pareto front: =$\mathcal{F}=\{V^{\pi}(s)\mid\pi\in\mathcal{P}\}.$
Since obtaining the true Pareto set is intractable in complex problems, the practical aim of multi-objective optimisation is to construct a finite set of policies that closely approximate the true Pareto front. So that practitioners can select the policy based on their preferred trade-off among objectives.
@@ 3 Methods
Figure 1: Parameter space and performance space of the Pareto policies. (Left) 2D projection of the high-dimensional policy parameter space. Red and blue gradient shadings and contour lines depict the scalarised reward under different preference vectors $\omega_{1}$ and $\omega_{2}$. The arrow marks a short retraining update. (Right) The policy obtained by retraining the $\theta_{w_{1}}$ model under $\omega_{2}$ (green) shifts towards the new preference as seen in performance space.
3.1 Overview
As shown in Fig. 1, we first explore the relationship between the parameter space and the performance space. We empirically find that a short retraining of a converged policy under a new preference induces a small, structured update in parameter space that corresponds to a predictable shift of the expected returns of the policy. This “model similarity”—the fact that the retrained policy stays close to the original parameters while already moving toward a different region of the Pareto front-–-underpins our method for steering policies along the front. Building on this insight, we explore the possibility that using the parameter-space difference between two structurally similar policies—trained under different preferences—to guide directional updates that extend our approximation of the Pareto front.
Leveraging this property, we develop an efficient algorithm to approximate the Pareto set of policies. We start by initializing a small collection of base policies, each trained to converge under a distinct scalarization weight chosen to span the preference evenly. Next, for each base policy, we perform a short retraining under a different preference weight, capturing the small parameter update that shifts the policy toward a new trade-off. These updates serve as directional moving vectors: we move from each base policy along its vector by a tunable step size to generate intermediate policies. Finally, we apply a brief fine-tuning to each intermediate policy under its corresponding preference, i.e. the scalarization weight shifted by the same fraction as the parameter updated, nudging it onto the true Pareto front.
@@ 3.2 Parameter-Performance Relationship
Recent work in multi-objective reinforcement learning has implicitly suggested a relationship between the parameter space of the policy network and the Pareto front in the performance space.  (Xu et al., 2020) empirically show for PGMORL that each disjoint policy family occupies a continuous region in parameter space and maps to a contiguous segment of the Pareto front, while MORL/D  (Felten et al., 2024) assume that policies with similar parameters should lead to close evaluations. Motivated by these implicit observations, we introduce the a parameter–performance relationship and proceed to explain and empirically validate this property.
Definition 3
Parameter-Performance Relationship (PPR). Let $\Theta\subseteq\mathbb{R}^{n}$ be the policy parameter space and $V:\Theta\to\mathbb{R}^{d}$ the mapping from parameter vectors $\theta$ to the expected return vectors $V(\theta)$. We say $V$ exhibits a continuous parameter–performance relationship on a region $U\subseteq\Theta$ if there exists a function $h:\mathbb{R}^{n}\to\mathbb{R}^{d}$ and a radius $\delta>0$ such that, for any $\theta\in U$ and any parameter perturbation $\Delta\theta$ with $\|\Delta\theta\|<\delta$ and $\theta+\Delta\theta\in U$, $V(\theta+\Delta\theta)\;-\;V(\theta)\;=\;h\bigl{(}\Delta\theta\bigr{)}.$
To study this relationship, we first need a metric for policy closeness in parameter space. We adopt the Hungarian matching distance (Kuhn, 1955; Munkres, 1957) to measure model distance and thereby quantify structural similarity between policies. Our policies are represented by feedforward neural networks with multiple layers. For each layer, we view the neurons of the two networks as two point-sets and compute a minimum-cost perfect matching between them, where the cost of matching neuron $i$ to neuron $j$ is the $\ell_{2}$ norm of the difference between their incoming weight vectors. Summing these layerwise minimal costs yields the total distance. This metric naturally handles the permutation invariance of hidden units (Goodfellow et al., 2016) and measures the smallest structural change needed to align one model to another—lower Hungarian distance indicates greater model similarity.
@@ 3.3 Sanity Check
To get a first idea about the PPR, we compare policies trained independently with those obtained by short retraining. We first train two policies to convergence using a multi-objective PPO-based (Schulman et al., 2017) algorithm with scalarization vectors $w_{1}$ and $w_{2}$, yielding model parameters $\theta_{w_{1}}$ and $\theta_{w_{2}}$. Starting from $\theta_{w_{1}}$, we then perform one short additional training step with $w_{2}$ to obtain $\theta_{w^{\prime}}$. To quantify how “close” these policy variants are, we show neuron heatmaps for each model both at the policy-network and value-network level in Figure 2(a), and visualise the Hungarian matching distances between those models in Figure 2(b). We also plot the rewards for three policies in the two-objective performance space (Figure 2(c)) for the multi-objective Swimmer problem.
We compare three pairs of models: (1) $\theta_{w_{1}}$ and $\theta_{w_{2}}$, capturing differences between independently trained policies in both parameter space and performance space; (2) $\theta_{w_{1}}$ and $\theta_{w^{\prime}}$, showing that brief retraining yields a structurally similar model and a low Hungarian matching distance, yet already shifted toward $w_{2}$ in reward space; and (3) $\theta_{w^{\prime}}$ and $\theta_{w_{2}}$, illustrating that although their parameters remain distinct, their rewards lie much closer on the performance space.
These empirical observations show that a short retraining step under a new preference produces a small, structured parameter update that directly maps to a predictable shift in performance, validating the local PPR.
(a) Policy-net and value-net parameter heatmap to illustrate the retraining effect that is measured in (b).
@@
Figure 2: Comparing independently trained policy $\theta_{w_{2}}$ versus retrained policy $\theta_{w^{\prime}}$ based on $\theta_{w_{1}}$, for details see Sect. 3.3. The environment used here is the multi-objective Swimmer problem.
@@ 3.4 Locally Linear Extension
Based on the PPR definition, a natural question is whether the parameter-space difference between two structurally similar policies—trained under different preferences—can serve as a directional update to extend our approximate Pareto front. To explore this, we consider two policies, a base policy $\theta_{w}$ and a retrained policy $\theta_{w^{\prime}}$, which exhibits a parameter-performance relationship. Crucially, for this directional information to be meaningful for Pareto front exploration, both $\theta_{w}$ and $\theta_{w^{\prime}}$ should ideally be non-dominated solutions, at least with respect to each other. Given such a pair, we compute the parameter update vector $\Delta\theta=\theta_{w^{\prime}}-\theta_{w}$ and generate a set of intermediate policies by moving from the base policy $\theta_{w}$ along the parameter displacement $\Delta\theta$ in scaled steps. Concretely, for each scale $\alpha$, we form $\theta_{\alpha}=\theta_{w}+\alpha\,\Delta\theta$  and evaluate its reward vectors in preference space.
Figure 3 visualises the resulting trajectory of reward vectors in the two-dimensional objective space: as $\alpha$ grows, the trajectory passes through the region around $\theta_{w^{\prime}}$ and can extend beyond both the base and retrained endpoints, demonstrating how simple parameter-space moves can traverse broad trade-off regions, which offers a cost-effective strategy for efficiently expanding an approximate Pareto front without training each point from scratch.
Figure 3: Visualisation of the process of applying the parameter difference $\Delta\theta=\theta_{w^{\prime}}-\theta_{w}$ between two related policies. The policies are obtained by first training a policy $\theta_{w}$ to convergence using scalarization vector $w$ and then find policy $\theta_{w^{\prime}}$ by a brief additional training period with a different scalarization vector $w^{\prime}$. Iterating the shift $\Delta\theta$ in the policy space induces a sequence of shifts also in the multi-objective reward space. The subfigures show results for different initial preferences: (a) A convex front is found from the two policies. (b) Although the original policy turns out to be Pareto suboptimal, the solution manifold extends into a Pareto optimal component. (c) Retraining can cause the (Pareto-suboptimal) original solution to jump to a different branch so that the corresponding solution consists of two components one of which can be ignored because of Pareto suboptimality.

### v2

@@ Abstract
Multi-objective reinforcement learning (MORL) aims at optimising several, often conflicting goals to improve the flexibility and reliability of RL in practical tasks. This is typically achieved by finding a set of diverse, non-dominated policies that form a Pareto front in the performance space. We introduce LLE-MORL, an approach that achieves interpretability by design by utilising a training scheme based on the local relationship between the parameter space and the performance space. By exploiting a locally linear map between these spaces, our method provides an interpretation of policy parameters in terms of the objectives, and this structured representation enables an efficient search within contiguous solution domains, allowing for the rapid generation of high-quality solutions without extensive retraining. Experiments across diverse continuous control domains demonstrate that LLE-MORL consistently achieves higher Pareto front quality and efficiency than state-of-the-art approaches.
Machine Learning, ICML
@@ 1 Introduction
While a globally continuous mapping is an ideal, we notice that in non-trivial problems, the relationship between the performance space and parameter space of policies is not a simple, single continuous mapping but can be described by a family of locally continuous components (Xu et al., 2020; Li et al., 2024). Our findings suggest that effectively exploring just a few of these components can be sufficient to achieve competitive performance in typical benchmark problems. To explore these components, we take a different perspective. Rather than searching for Pareto-optimal policies independently, we aim to trace the Pareto front manifold directly by exploring local structure in the high-dimensional policy parameter space. We formalise the Parameter-Performance Relationship (PPR), which captures the local correspondence between policy parameters and objective performance. This viewpoint enables an efficient strategy for Pareto front (see Sect. 2.2) approximation: starting from a small set of policies, we can expand along directions that preserve PPR and generate new candidate solutions via a training-free extension stage, establishing an explicit and interpretable relationship between the policy parameter space and the reward space.
Based on this perspective, we develop a novel algorithm LLE-MORL (Locally Linear Extrapolation for Multi-Objective Reinforcement Learning). Our experiments demonstrate that the proposed algorithm can achieve high-quality Pareto front approximations with notable sample efficiency compared to existing approaches. By focusing on local structure in the policy parameter space, LLE-MORL reduces the need for exhaustive training while maintaining strong coverage and solution quality. Beyond efficiency, the structured use of the PPR provides an interpretable representation of the policy set, supporting a clearer understanding of local trade-offs along the Pareto front.
@@ 2.1 Multi-Objective Reinforcement Learning
Definition 2.1.
Multi-Objective Markov Decision Process (MOMDP). A MOMDP is defined by the tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the state transition probability, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is a vector-valued reward function with $d$ as the number of objectives, specifying the immediate reward for each of the objectives, $\gamma$ is the discount factor.
The crucial difference between MOMDPs and traditional single-objective MDPs is the reward structure. While single-objective MDPs use a scalar reward function, MOMDPs feature a vector-valued reward function that outputs $d$-dim vector rewards $\mathbf{r}=R(s,a)$, which delivers distinct numeric feedback for each objective. We denote a policy by $\pi$. At each timestep $t$, the agent in state $s_{t}\in\mathcal{S}$ selects an action $a_{t}\sim\pi(\cdot\mid s_{t})$, transitions to a new state $s_{t+1}$ with probability $P(s_{t+1}\mid s_{t},a_{t})$, and receives a reward vector $\mathbf{r}_{t}=\bigl[(R_{1}(s_{t},a_{t}),\,R_{2}(s_{t},a_{t}),\,\dots,\,R_{d}(s_{t},a_{t})\bigr]).$ We define the discounted return vector by $\mathbf{G}_{t}=\sum_{k=0}^{\infty}\gamma^{k}\,\mathbf{r}_{t+k},$ and the multi-objective action-value function of a policy $\pi$ for a given state-action pair $(s,a)$ by $\mathbf{Q}^{\pi}(s,a)=\mathbb{E}_{\pi}\bigl[\mathbf{G}_{t}\mid s_{t}=s,\,a_{t}=a\bigr]$. The goal of MORL is to find a policy $\pi$ such that the expected return of each objective can be optimised. In practice, we trade off objectives via a scalarisation function $f_{\boldsymbol{\omega}}(\mathbf{r})$, which produces a scalar utility using preference vector $\boldsymbol{\omega}\in\Omega$, where $\Omega$ is the preferences space. The scalarisation function $f_{\boldsymbol{\omega}}(\mathbf{r})$ is used for mapping the multi-objective reward vector $\mathbf{r}(s,a)$ to a single scalar. In this paper, we consider the linear scalarisation function $f_{\boldsymbol{\omega}}(\mathbf{r}(s,a))=\boldsymbol{\omega}^{\mathbf{T}}\mathbf{r}(s,a)$, where the preference vector $\boldsymbol{\omega}$ lies in the probability simplex, i.e., $\sum_{i=1}^{d}\omega_{i}=1$ and $\omega_{i}\geq 0$, which is commonly used in MORL literature (Yang et al., 2019; Felten et al., 2024). When the preference dimension $d=1$, the MOMDP reduces to a standard single-objective MDP under this linear scalarisation setting.
@@ 2.2 Pareto Optimality
Definition 2.2.
Pareto optimality. A policy $\pi$ is said to dominate another policy $\pi^{\prime}$ if and only if $\forall i\in\{1,\dots,d\},\quad V_{i}^{\pi}(s)\geq V_{i}^{\pi^{\prime}}(s),\quad\text{and}\quad\exists j,\,V_{j}^{\pi}(s)>V_{j}^{\pi^{\prime}}(s),$ where $V_{i}^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{i}(s_{t},a_{t})\mid s_{0}=s]$ denotes the expected discounted return for objective $i$ under policy $\pi$. A policy $\pi^{*}$ is Pareto optimal if and only if it is not dominated by another policy. The set of all Pareto optimal policies forms the Pareto set: $\Pi_{P}=\{\pi\mid\pi\text{ is Pareto optimal}\}.$ The corresponding set of expected returns incured by policies in the Pareto set is termed Pareto front: $\mathcal{P}=\{V^{\pi}(s)\mid\pi\in\Pi_{P}\}.$
Since obtaining the true Pareto set is intractable in complex problems, the practical aim of multi‐objective optimisation is to construct a finite set of policies that closely approximate the true Pareto front. So that practitioners can select the policy based on their preferred trade-off among objectives.
@@ 3 Methods
Figure 1: Parameter space and performance space of the Pareto policies. (Left) 2D projection of the high‑dimensional policy parameter space. Red and blue gradient shadings and contour lines depict the scalarised reward under different preference vectors $\omega_{1}$ and $\omega_{2}$. The arrow marks a short retraining update. (Right) The policy obtained by retraining the $\theta_{w_{1}}$ model under $\omega_{2}$ (green) shifts towards the new preference as seen in performance space.
@@ 3.1 Parameter-Performance Relationship
Recent work in MORL has implicitly suggested a relationship between the parameter space of the policy network and the Pareto front in the performance space.  (Xu et al., 2020) empirically show for PGMORL that each disjoint policy family occupies a continuous region in parameter space and maps to a contiguous segment of the Pareto front, while MORL/D  (Felten et al., 2024) assume that policies with similar parameters should lead to close evaluations. Motivated by these implicit observations, we introduce the a parameter–performance relationship and proceed to explain and empirically validate this property.
Definition 3.1.
Parameter-Performance Relationship (PPR). Let $\Theta\subseteq\mathbb{R}^{n}$ be the space of policy parameters, and $V$ be the expected discounted return as a function $V:\Theta\to\mathbb{R}^{d}$ of the parameter $\theta$. We say that $V$ exhibits a parameter–performance relationship on an open region $U\subseteq\Theta$ if $\exists\delta>0$ and a function $h:U\times B_{\delta}(0)\to\mathbb{R}^{d}$, s.t. $\forall\theta\in U$ and perturbation $\Delta\theta$ satisfying $\|\Delta\theta\|<\delta$ and $\theta+\Delta\theta\in U$, the difference in the performance space is:
$V(\theta+\Delta\theta)\;-\;V(\theta)\;=\;h\bigl(\theta,\Delta\theta\bigr).$ (1)
where $B_{\delta}(0)$ is an open ball with radius $\delta$ centres on the origin, i.e. $B_{\delta}(0):=\{x:\|x\|<\delta\}$.
PPR is a local property and need not hold globally across the entire parameter space. In high-dimensional parameter spaces, policies with similar performance may be far apart in parameter space. This raises the question of how to identify policies that satisfy the PPR locally. As a first step, we require a metric for policy closeness in parameter space. We adopt the Hungarian matching distance (Kuhn, 1955; Munkres, 1957) to measure model distance and thereby quantify structural similarity between policies. A formal definition of Hungarian matching distance is provided in Appendix C.3. This metric naturally handles the permutation invariance of hidden units (Goodfellow et al., 2016) and measures the smallest structural change needed to align one model to another: lower Hungarian distance indicates greater model similarity.
As illustrated in Figure 1, a short retraining of a policy under a nearby preference typically induces a small, structured update in parameter space that corresponds to a predictable shift in the expected returns. This observation motivates our use of brief retraining as a practical means of obtaining locally related policies for which the PPR holds.
@@ 3.2 Sanity Check
To obtain an initial empirical understanding of the PPR, we compare independently trained policies with those obtained via brief retraining. We first train two policies to a stable stage using a multi-objective PPO-based (Schulman et al., 2017) algorithm with scalarization vectors $w_{1}$ and $w_{2}$, yielding model parameters $\theta_{w_{1}}$ and $\theta_{w_{2}}$. Starting from $\theta_{w_{1}}$, we then perform one short additional training step with $w_{2}$ to obtain $\theta_{w^{\prime}}$. We qualitatively evaluated the similarity between these policies using neuron heat maps at both the policy and value network levels in Figure 2(a), and quantitatively by computing the Hungarian matching distances between model pairs in Figure 2(b). Their corresponding rewards are visualised in the two-objective performance space for the multi-objective Swimmer problem in Figure 2(c).
We compare three pairs of models: (1) $\theta_{w_{1}}$ and $\theta_{w_{2}}$, representing independently trained policies that differ substantially in both parameter and performance space; (2) $\theta_{w_{1}}$ and $\theta_{w^{\prime}}$, showing that brief retraining produces a structurally similar policy with a low Hungarian distance while inducing a clear shift in performance; and (3) $\theta_{w^{\prime}}$ and $\theta_{w_{2}}$, demonstrating that although their parameters remain distinct, their rewards lie much closer in the performance space.
These empirical observations indicate that a short retraining under a new preference produces a small, structured parameter update that directly maps to a predictable shift in performance, validating the PPR. More details of the sanity check procedure are provided in Appendix C.4.
(a) Policy‐net and value-net parameter heatmap. The horizontal axis corresponds to network layers, the vertical axis corresponds to neurons within each layer, and colour indicates the corresponding parameter values.
@@
Figure 2: Comparing independently trained policy $\theta_{w_{2}}$ versus retrained policy $\theta_{w^{\prime}}$ based on $\theta_{w_{1}}$, for details see Section 3.2. The environment used here is the multi-objective Swimmer problem.
@@ 3.3 Locally Linear Extension
Based on the PPR definition, a natural question is whether the parameter-space difference between two policies that satisfy the PPR can serve as a directional update to extend an approximate Pareto front. To explore this, we consider two policies, a base policy $\theta_{w}$ and a retrained policy $\theta_{w^{\prime}}$, which exhibit a PPR. Crucially, for this directional information to be meaningful for Pareto front exploration, both $\theta_{w}$ and $\theta_{w^{\prime}}$ should ideally be non-dominated solutions, at least with respect to each other. Given such a pair, we compute the parameter update vector $\Delta\theta=\theta_{w^{\prime}}-\theta_{w}$ and generate a set of intermediate policies by moving from the base policy $\theta_{w}$ along the parameter displacement $\Delta\theta$ in scaled steps. Concretely, for each scale $\alpha$, we form $\theta_{\alpha}=\theta_{w}+\alpha\,\Delta\theta$  and evaluate its reward vectors in preference space.
Figure 3 visualises the resulting trajectory of reward vectors in the two-dimensional objective space: as $\alpha$ grows, the trajectory passes through the region around $\theta_{w^{\prime}}$ and can extend beyond both the base and retrained endpoints, demonstrating how simple parameter‐space moves can traverse broad trade‑off regions, which offers a cost‑effective strategy for efficiently expanding an approximate Pareto front without training each point from scratch.
Figure 3: Visualisation of the process of applying the parameter difference $\Delta\theta=\theta_{w^{\prime}}-\theta_{w}$ between two related policies. The policies are obtained by first training a policy $\theta_{w}$ to a stable stage using scalarization vector $w$ and then find policy $\theta_{w^{\prime}}$ by a brief additional training period with a different scalarization vector $w^{\prime}$. Iterating the shift $\Delta\theta$ in the policy space induces a sequence of shifts also in the multi-objective reward space. The subfigures show results for different initial preferences: (a) A convex front is found from the two policies. (b) Although the original policy turns out to be Pareto suboptimal, the solution manifold extends into a Pareto optimal component. (c) Retraining can cause the (Pareto-suboptimal) original solution to jump to a different branch so that the corresponding solution consists of two components one of which can be ignored because of Pareto suboptimality.